OpenStudy (anonymous):

]Choose the equation below that represents the line passing through the point (2, -5) with a slope of -3.

OpenStudy (falco276):

Is this supposed to be Algebra?

OpenStudy (anonymous):

Yessa

OpenStudy (falco276):

I really don't want to be mean to you, but are you the real Chief Keef? I saw one of your vids on Youtube. Youre awesome. :D

OpenStudy (anonymous):

haha yessa

OpenStudy (falco276):

I mean really? Wow. Keep going with your songs, theyre really good! ahem.... back to Algebra, I'm horrible. :P

OpenStudy (anonymous):

#300

OpenStudy (falco276):

????

OpenStudy (anonymous):

are you gonna help me ?

OpenStudy (falco276):

I might not cuz I suck at Algebra..... Here, I'll tag you in a few experts, hopefully they'll arrive. @primeralph @genius12 @nincompoop @Preetha Hope that helps. :)

OpenStudy (primeralph):

@AlmightySosa300 Where are the "equations below"?

OpenStudy (anonymous):

y = -3x - 13 y = -3x + 11 y = -3x + 13 y = -3x + 1

OpenStudy (primeralph):